# Optimal Projection for 3D Gaussian Splatting

### v2

@@ 4 Local Affine Approximation Error
From Equation 6 and Equation 7 we observe that 3D Gaussian Splatting introduces an approximation during the projection transformation, i.e. the local affine approximation. The rasterization uses a 2D Gaussian function that is not the true projection function of the 3D Gaussian. The error introduced by this approximation can be characterized by the Taylor remainder term in Equation 6:
$\begin{split}&R_{1}\left(\mathbf{x}^{{}^{\prime}}\right)=\phi\left(\mathbf{x}^{{}^{\prime}}\right)-\phi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)-\frac{\partial\phi}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\left(\mathbf{x}^{{}^{\prime}}-\boldsymbol{\mu}^{{}^{\prime}}\right)\end{split}$ (9)
@@
$\frac{\partial\phi}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)=\mathbb{I}\otimes\left(\mathbf{x_{0}}^{T}\boldsymbol{\mu}^{{}^{\prime}}\right)^{-1}-\mathbf{x_{0}}\left(\mathbf{x_{0}}^{T}\boldsymbol{\mu}^{{}^{\prime}}\right)^{-1}\left({\boldsymbol{\mu}^{{}^{\prime}}}^{T}\mathbf{x_{0}}\right)^{-1}{\boldsymbol{\mu}^{{}^{\prime}}}^{T}$ (10)
and $\mathbb{I}$ represents the identity matrix; $\otimes$ denotes the multiplication of a matrix and a scalar. It is not difficult to observe that this Taylor remainder term is a three-dimensional vector related to the random variable $\mathbf{x}^{{}^{\prime}}$ and the mean $\boldsymbol{\mu}^{{}^{\prime}}$ of the Gaussian function. Therefore, by computing the square of the Frobenius norm of this vector $\|R_{1}\left(\mathbf{x}^{{}^{\prime}}\right)\|_{F}^{2}$ and taking the mathematical expectation of this norm function with respect to the random variable $\mathbf{x}^{{}^{\prime}}$, we ultimately obtain an error function that depends solely on the Gaussian mean $\boldsymbol{\mu}^{{}^{\prime}}$:
@@
Before seeking the extremum of the error function, let’s simplify the integral expression. First, we simplify $\mathbf{x}^{{}^{\prime}}$ and $\boldsymbol{\mu}^{{}^{\prime}}$ to unit vectors projected onto the unit sphere centered at the camera center (hereafter referred to as the unit sphere) by proving that the composition of transformations $\left(\phi\circ\pi\right)$, projecting a point in three-dimensional space onto the unit sphere $\pi$ and subsequently onto the projection plane $\phi$, is equivalent to directly projecting a point in three-dimensional space onto the projection plane. This simplification ensures that the entire error function involves only operations between three unit space vectors. The proof proceeds as follows:
$\begin{split}\left(\phi\circ\pi\right)\left({\mathbf{x}^{{}^{\prime}}}\right)&=\phi\left({\mathbf{x}^{{}^{\prime}}}{\left({\mathbf{x}^{{}^{\prime}}}^{\top}{\mathbf{x}^{{}^{\prime}}}\right)^{-1/2}}\right)\\ &={\mathbf{x}^{{}^{\prime}}}{\left({\mathbf{x}^{{}^{\prime}}}^{\top}{\mathbf{x}^{{}^{\prime}}}\right)^{-1/2}}\left(\mathbf{x_{0}}^{\top}\left({\mathbf{x}^{{}^{\prime}}}{\left({\mathbf{x}^{{}^{\prime}}}^{\top}{\mathbf{x}^{{}^{\prime}}}\right)^{-1/2}}\right)\right)^{-1}\\ &={\mathbf{x}^{{}^{\prime}}}\left(\mathbf{x_{0}}^{T}{\mathbf{x}^{{}^{\prime}}}\right)^{-1}\\ &=\phi\left({\mathbf{x}^{{}^{\prime}}}\right)\text{.}\end{split}$ (12)
The simplified error function $\epsilon$ involves three unit vectors, or equivalently, three points on the unit sphere: $\mathbf{x}^{{}^{\prime}}$, $\mathbf{x_{0}}$ and $\boldsymbol{\mu}^{{}^{\prime}}$, as illustrated in Figure 3. Their spherical coordinates are given as follows:
@@
where $\phi_{0}=0$ and $\theta_{0}=0$. Based on Equation 9, Equation 10, and Equation 13, the following derivations can be obtained:
$\mathbf{J}=\frac{\partial\phi}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)=\left[\begin{matrix}\frac{1}{\cos{\left(\phi_{\mu}\right)}\cos{\left(\theta_{\mu}\right)}}&0&-\frac{\sin{\left(\phi_{\mu}\right)}}{\cos^{2}{\left(\phi_{\mu}\right)}\cos{\left(\theta_{\mu}\right)}}\\ 0&\frac{1}{\cos{\left(\phi_{\mu}\right)}\cos{\left(\theta_{\mu}\right)}}&\frac{\sin{\left(\theta_{\mu}\right)}}{\cos^{2}{\left(\phi_{\mu}\right)}\cos^{2}{\left(\theta_{\mu}\right)}}\\ 0&0&0\end{matrix}\right]$ (14)
The size of the integration region is related to the covariance of the Gaussian function. Since the current discussion specifically focuses on the impact of the mean on the error rather than the variance, let’s assume the integration region to be $\left\{\mathbf{x}^{{}^{\prime}}\mid\theta\in\left[-\pi/4+\theta_{\mu},\pi/4+\theta_{\mu}\right]~\land~\phi\in\left[-\pi/4+\phi_{\mu},\pi/4+\phi_{\mu}\right]\right\}$. Then, substituting Equation 15 into Equation 11, we can derive:
@@
The function is partially differentiated with respect to $\theta_{\mu}$ and $\phi_{\mu}$. It is observed that at $\theta_{\mu}=\theta_{0}=0$ and $\phi_{\mu}=\phi_{0}=0$, the following holds:
@@
Therefore, this point is an extremum point of the function. Upon verification, it is confirmed that this point is the minimum value point of the function. We have visualized the error function, as shown in Figure 2. Through Subfigure 2a in Figure 2, it can be observed that this function is a concave function, and it attains its minimum value at (0,0)(0,0)( 0 , 0 ). Examining Subfigure 2b in Figure 2, it is evident that in the majority of the region close to the origin, the function values are not significantly different. However, as the function approaches the integration limits, the function values increase rapidly, resulting in a substantial difference between the maximum and minimum values.
In other words, in most cases, the error is small and does not significantly affect the quality of rendered images. This is why 3D Gaussian Splatting utilizes local affine approximation but still successfully reconstructs the scene to obtain a high-quality novel view image. Nevertheless, the naive projection of all gaussians onto the same plane $z=1$ in 3D Gaussian Splatting may lead to larger projection errors for gaussians farther from the plane center (the projection point of the camera center to the plane), causing artifacts, as shown in Figure 5.
Figure 3: Toy 2D visualizations of the projection transformation. $\mathbf{x}^{{}^{\prime}}$, $\mathbf{x_{0}}$ and $\boldsymbol{\mu}^{{}^{\prime}}$ are all simplified unit vectors, representing the random variable, the tangency point of the projection plane and the unit sphere, and the Gaussian mean, respectively, denoted by green, purple, and blue arrows in the diagram. Due to the $3\sigma$ rule in the Gaussian distribution and the fact that actual rasterization does not sample all points, we represent the integration domain with a circular arc symmetric about the mean vector $\boldsymbol{\mu}^{{}^{\prime}}$.
@@ 5 Optimal Gaussian Splatting
Figure 4: Illustration of the rendering pipeline for Optimal Gaussian Splatting. The first step, similar to the original 3D-GS [16], involves transforming world coordinates into the camera coordinate system. Instead of projecting Gaussians onto the z=1 plane, each Gaussian is radially projected based on its mean along the line connecting it to the camera center, projecting onto a plane tangent to the unit sphere and perpendicular to this line. After projection, colors of points on the unit sphere are computed by alpha blending the 2D Gaussians on the tangent plane. Finally, for each pixel in the image, rays are cast onto the unit sphere to retrieve the color of that pixel, resulting in the rendered image.
Based on the analysis of the error function, in this section, we will introduce Optimal Gaussian Splatting shown in Figure 4 that exhibits smaller projection errors, resulting in higher-quality rendering, as shown in Figure 5 and Table 1.
@@ 5.1 Optimal Projection
Through the analysis of the error function in Section 4, we discovered that the error function attains its minimum value when the projection of the Gaussian mean on the plane coincides with the projection from the camera center to the plane. Therefore, in Optimal Gaussian Splatting, we employ an optimal projection method. Specifically, instead of naively projecting different Gaussians onto the same plane $z=1$, we adopt distinct projection planes for each Gaussian. These projection planes are determined by the tangent planes formed by the Gaussian mean and the line connecting it to the camera center. The specific projection plane is ormulated as:
$\mathbf{x^{\mu}}^{\top}\cdot\left(\mathbf{x}^{{}^{\prime}}-\mathbf{x^{\mu}}\right)=0$ (18)
@@
$\mathbf{x^{\mu}}=\pi\left({\boldsymbol{\mu}^{{}^{\prime}}}\right)={\boldsymbol{\mu}^{{}^{\prime}}}\left({\boldsymbol{\mu}^{{}^{\prime}}}^{\top}{\boldsymbol{\mu}^{{}^{\prime}}}\right)^{-1/2}$ (19)
represents the projection of the camera coordinate system origin onto this plane. According to the projection plane equation (Equation 18 and Equation19), the optimal projection function $\phi_{o}$ is obtained as:
$\phi^{\mu}\left(\mathbf{x}^{{}^{\prime}}\right)=\mathbf{x}^{{}^{\prime}}\left(\mathbf{x^{\mu}}^{T}\mathbf{x}^{{}^{\prime}}\right)^{-1}\left(\mathbf{x^{\mu}}^{T}\mathbf{x^{\mu}}\right)=\mathbf{x}^{{}^{\prime}}\left(\mathbf{x^{\mu}}^{T}\mathbf{x}^{{}^{\prime}}\right)^{-1}\text{.}$ (20)
The corresponding local affine approximation Jacobian matrix $\mathbf{J_{o}}$ is:
$\mathbf{J^{\mu}}=\frac{\partial\phi^{\mu}}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)=\mathbb{I}\otimes\left(\mathbf{x^{\mu}}^{T}\boldsymbol{\mu}^{{}^{\prime}}\right)^{-1}-\mathbf{x^{\mu}}\left(\mathbf{x^{\mu}}^{T}\boldsymbol{\mu}^{{}^{\prime}}\right)^{-1}\left({\boldsymbol{\mu}^{{}^{\prime}}}^{T}\mathbf{x^{\mu}}\right)^{-1}{\boldsymbol{\mu}^{{}^{\prime}}}^{T}\text{.}$ (21)
@@
$\mathbf{J^{\mu}}=\left[\begin{matrix}\frac{\mu_{y}^{2}+\mu_{z}^{2}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}&-\frac{\mu_{x}\mu_{y}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}&-\frac{\mu_{x}\mu_{z}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}\\ -\frac{\mu_{x}\mu_{y}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}&\frac{\mu_{x}^{2}+\mu_{z}^{2}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}&-\frac{\mu_{y}\mu_{z}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}\\ -\frac{\mu_{x}\mu_{z}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}&-\frac{\mu_{y}\mu_{z}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}&\frac{\mu_{x}^{2}+\mu_{y}^{2}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}\end{matrix}\right]$ (22)
@@ 5.2 Unit Sphere Based Rasterizer
For Optimal Projection, we have employed a Unit Sphere Based Rasterizer for rasterization to generate images.
@@
Through Optimal Projection, we obtain the projection of the three-dimensional Gaussians on the tangent plane of the unit sphere instead of obtaining the Gaussians in the image space. Therefore, we need to rasterize based on this unit sphere to generate images.
@@
In specific terms, for a pixel $\left(u,v\right)$ on the image, similar to NeRF, we cast a ray. However, unlike NeRF, we do not involve the extensive sampling of points along the ray, which would significantly degrade performance. Instead, our focus is solely on determining which tangent-plane Gaussians the ray intersects on the unit sphere:
$\mathbf{x^{\mu}_{\text{2D}}}=\phi^{\mu}\left(\left[\begin{matrix}\left(u-c_{x}\right)/f_{x}\\ \left(v-c_{y}\right)/f_{y}\\ 1\end{matrix}\right]\right)$ (23)
@@
We then query the function values of these Gaussians for alpha blending to obtain the color
@@
$G^{\mu}_{\text{2D}}\left(\mathbf{x^{\mu}_{\text{2D}}}\right)=\exp\left(-1/2\left(\mathbf{x^{\mu}_{\text{2D}}}-\phi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\right)^{\top}\left(\mathbf{J^{\mu}}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J^{\mu}}^{\top}\right)^{-1}\left(\mathbf{x^{\mu}_{\text{2D}}}-\phi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\right)\right)$ (24)
$\alpha\left(u,v,\boldsymbol{\mu}\right)=\alpha_{\boldsymbol{\mu}}\cdot G^{\mu}_{\text{2D}}\left(\mathbf{x^{\mu}_{\text{2D}}}\right)$ (25)
$\mathbf{C}\left(u,v\right)=\sum_{z=z_{\text{near}}}^{z_{\text{far}}}{\text{SH}\left(\boldsymbol{\mu}_{z}\right)\alpha\left(u,v,\boldsymbol{\mu}_{z}\right)\prod_{k=z_{\text{near}}}^{z}\left(1-\alpha\left(u,v,\boldsymbol{\mu}_{k}\right)\right)}$ (26)
where the summation and product from $z_{\text{near}}$ to $z_{\text{far}}$ represent discretized depths sorted in ascending order from near to far for the depths of 3D Gaussians encountered by the ray, $\alpha_{\boldsymbol{\mu}}$ represents the transparency of the 3D Gaussian with its mean $\boldsymbol{\mu}$, $\boldsymbol{\mu}_{z}$ represents the mean of the Gaussian at depth $z$, and SH represents the spherical harmonics of this Gaussian.
@@
It is easy to notice that both $\mathbf{J}^{\mu}$ in Equation 21 and $\mathbf{J}$ in Equation 10 are matrices with a rank of 2. Therefore, in fact, $\mathbf{J^{\mu}}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J^{\mu}}^{\top}$ does not have an inverse. We perform a reversible transformation on $\mathbf{J}^{\mu}$ to make the third row zero. Consequently, $\mathbf{J^{\mu}}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J^{\mu}}^{\top}$ becomes a matrix with both the third row and the third column being zero. This allows us to treat it similar to $\mathbf{J}$, skipping the third row and third column, and compute its inverse as if it were a full-rank $\times 2$ 2 matrix. Certainly, to ensure the equation still holds, this invertible matrix also needs to be used for left multiplication with $\mathbf{x^{\mu}_{\text{2D}}}$ and $\phi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)$.

### v4

@@ 4 Local Affine Approximation Error
From Equation 3 and Equation 4 we observe that 3D Gaussian Splatting introduces an approximation during the projection transformation, i.e. the local affine approximation. Consequently, 3D-GS actually adopts a 2D Gaussian function that is not the true projection of the initial 3D Gaussian. The error introduced by this approximation can be characterized by the Taylor remainder term in Equation 3:
$\begin{split}&R_{1}\left(\mathbf{x}^{{}^{\prime}}\right)=\varphi\left(\mathbf{x}^{{}^{\prime}}\right)-\varphi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)-\frac{\partial\varphi}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\left(\mathbf{x}^{{}^{\prime}}-\boldsymbol{\mu}^{{}^{\prime}}\right)\end{split}$ (5)
@@
$\frac{\partial\varphi}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)=\mathbb{I}\otimes\left(\mathbf{x_{0}}^{\top}\boldsymbol{\mu}^{{}^{\prime}}\right)^{-1}-{\boldsymbol{\mu}^{{}^{\prime}}}\left(\mathbf{x_{0}}^{\top}\boldsymbol{\mu}^{{}^{\prime}}\right)^{-1}\left({\boldsymbol{\mu}^{{}^{\prime}}}^{\top}\mathbf{x_{0}}\right)^{-1}\mathbf{x_{0}}^{\top}\text{.}$ (6)
$\mathbb{I}$ represents the identity matrix and $\otimes$ denotes the multiplication of a matrix and a scalar. Clearly, this Taylor remainder term is a three-dimensional vector related to the random variable $\mathbf{x}^{{}^{\prime}}$ and the mean $\boldsymbol{\mu}^{{}^{\prime}}$ of the Gaussian function. Therefore, by computing the square of the Frobenius norm of this vector $\|R_{1}\left(\mathbf{x}^{{}^{\prime}}\right)\|_{F}^{2}$ and taking the mathematical expectation of this norm function with respect to the random variable $\mathbf{x}^{{}^{\prime}}$, we obtain an error function that depends solely on the Gaussian mean $\boldsymbol{\mu}^{{}^{\prime}}$:
@@
The simplified error function $\epsilon$ involves three unit vectors: $\mathbf{x}^{{}^{\prime}}$, $\mathbf{x_{0}}$ and $\boldsymbol{\mu}^{{}^{\prime}}$. Their spherical coordinates are given as follows:
@@
where $\phi_{0}=0$ and $\theta_{0}=0$. Please refer to the detailed proofs in the supplementary materials.
The size of the integration region is related to the covariance of the Gaussian function. Since the current discussion specifically focuses on the impact of the mean on the error rather than the variance, let’s assume the integration region to be $\left\{\mathbf{x}^{{}^{\prime}}\mid\theta\in\left[-\pi/4+\theta_{\mu},\pi/4+\theta_{\mu}\right]~{}\land~{}\phi\in\left[-\pi/4+\phi_{\mu},\pi/4+\phi_{\mu}\right]\right\}$. Then, based on Equation 5-8, we obtain:
@@
Fortunately, this integral has a close-form expression, but its form is very complicated. We report the close-form expression in the supplemental materials and visually depict it in Figure 2.
Figure 2: Visualization of the 3D Gaussian Splatting error function $\epsilon\left(\theta_{\mu},\phi_{\mu}\right)$ under two distinct domains. The domain of the function is $\left\{\left(\theta_{\mu},\phi_{\mu}\right)\mid\theta_{\mu}\in\left[-\lambda\pi/4,\lambda\pi/4\right]~{}\land~{}\phi_{\mu}\in\left[-\lambda\pi/4,\lambda\pi/4\right]\right\}$ where $\left(\theta_{\mu},\phi_{\mu}\right)$ represents the polar coordinate of the 3D Gaussian mean and $\lambda$ represents the scaling factor of the domain.
The error function is partially differentiated with respect to $\theta_{\mu}$ and $\phi_{\mu}$. It is easy to check that at $\theta_{\mu}=\theta_{0}=0$ and $\phi_{\mu}=\phi_{0}=0$, the following holds:
@@
Therefore, this point is actually an extremum point of the function. We also confirm that this point is the minimum of the function. Indeed, from Figure 2(a), it is also apparent that the function is concave, with its minimum value occurring at $(0,0)$. Additionally, it can be observed that the minimum value of the error function is greater than zero, due to the fact that Gaussian functions are not closed under projection transformations. Figure 2(b) shows that in the majority of the region close to the origin, the function values vary very gently. However, as the independent variable of the function approaches the boundaries, the function values increase rapidly, resulting in a substantial difference between the maximum and minimum values.
The above analysis indicates that the error is small and does not significantly affect the quality of rendered images. Therefore, traditional 3D-GS [17] utilizes local affine approximation but still successfully reconstructs the scene to obtain a high-quality novel view image. Nevertheless, the naive projection of all Gaussians onto the same plane $z=1$ in 3D-GS may lead to larger projection errors for Gaussians farther from the plane center (the projection point of the camera center to the plane), causing artifacts. These artifacts are severe when using wide-angle lenses, significantly degrading the quality of the renderings.
@@ 5 Optimal Gaussian Splatting
Figure 3: Illustration of the rendering pipeline for our Optimal Gaussian Splatting and the projection of 3D-GS [17]. The blue box depicts the projection process of the original 3D-GS, which straightforwardly projects all Gaussians onto the same projection plane. In contrast, the red box illustrates our approach, where we project individual Gaussians onto corresponding tangent planes.
Based on the analysis of the error function, we introduce Optimal Gaussian Splatting shown in Figure 3. This helps us to achieve smaller projection errors, resulting in higher-quality rendering. The first step, similar to the original 3D-GS [17], involves transforming world coordinates into the camera coordinate system. Instead of projecting Gaussians onto the $z=1$ plane, each Gaussian is radially projected based on its mean along the line connecting it to the camera center, projecting onto a plane tangent to the unit sphere and perpendicular to this line. After projection, colors of points on the unit sphere are computed by alpha blending the 2D Gaussians on the tangent plane. Finally, for each pixel in the image, rays are cast onto the unit sphere to retrieve the color of that pixel, resulting in the rendered image.
@@ 5.1 Optimal Projection
We discover that the error function attains its minimum value when the projection of the Gaussian mean on the plane coincides with the projection from the camera center to the plane. Therefore, in Optimal Gaussian Splatting, we employ an optimal projection method. Specifically, instead of naively projecting different Gaussians onto the same plane $z=1$, we adopt distinct projection planes for each Gaussian. These projection planes are determined by the tangent planes formed by the Gaussian mean and the line connecting it to the camera center. The specific projection plane is formulated as:
$\mathbf{x_{p}}^{\top}\cdot\left(\mathbf{x}^{{}^{\prime}}-\mathbf{x_{p}}\right)=0$ (11)
@@
$\mathbf{x_{p}}=\varpi\left({\boldsymbol{\mu}^{{}^{\prime}}}\right)={\boldsymbol{\mu}^{{}^{\prime}}}\left({\boldsymbol{\mu}^{{}^{\prime}}}^{\top}{\boldsymbol{\mu}^{{}^{\prime}}}\right)^{-1/2}$ (12)
represents the projection of the camera space’s origin onto this plane and $\varpi$ function represents the transformation that projects points onto the unit sphere. According to Equation 11, the optimal projection function $\varphi_{\text{p}}$ is obtained as:
$\varphi_{\text{p}}\left(\mathbf{x}^{{}^{\prime}}\right)=\mathbf{x}^{{}^{\prime}}\left(\mathbf{x_{p}}^{\top}\mathbf{x}^{{}^{\prime}}\right)^{-1}\left(\mathbf{x_{p}}^{\top}\mathbf{x_{p}}\right)=\mathbf{x}^{{}^{\prime}}\left(\mathbf{x_{p}}^{\top}\mathbf{x}^{{}^{\prime}}\right)^{-1}\text{.}$ (13)
The corresponding local affine approximation Jacobian matrix $\mathbf{J_{p}}$ is:
$\mathbf{J_{p}}=\frac{\partial\varphi_{\text{p}}}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)=\mathbb{I}\otimes\left(\mathbf{x_{p}}^{\top}\boldsymbol{\mu}^{{}^{\prime}}\right)^{-1}-{\boldsymbol{\mu}^{{}^{\prime}}}\left(\mathbf{x_{p}}^{\top}\boldsymbol{\mu}^{{}^{\prime}}\right)^{-1}\left({\boldsymbol{\mu}^{{}^{\prime}}}^{\top}\mathbf{x_{p}}\right)^{-1}\mathbf{x_{p}}^{\top}\text{.}$ (14)
@@
$\mathbf{J_{p}}=\left[\begin{matrix}\frac{\mu_{y}^{2}+\mu_{z}^{2}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}&-\frac{\mu_{x}\mu_{y}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}&-\frac{\mu_{x}\mu_{z}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}\\ -\frac{\mu_{x}\mu_{y}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}&\frac{\mu_{x}^{2}+\mu_{z}^{2}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}&-\frac{\mu_{y}\mu_{z}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}\\ -\frac{\mu_{x}\mu_{z}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}&-\frac{\mu_{y}\mu_{z}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}&\frac{\mu_{x}^{2}+\mu_{y}^{2}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}\end{matrix}\right]\text{.}$ (15)
It’s very easy to implement modifications to the Jacobian matrix in the forward process.
@@ 5.2 Unit Sphere Based Rasterizer
For the proposed Optimal Projection, we have proposed a Unit Sphere Based Rasterizer to generate images. Through Optimal Projection, we obtain the projection of the three-dimensional Gaussians on the tangent plane of the unit sphere instead of obtaining the Gaussians in the image space. Therefore, we need to rasterize based on this unit sphere.
@@
Specifically, for a pixel $\left(u,v\right)$ on the image, similar to NeRF, we cast a ray. However, unlike NeRF, we do not involve the extensive sampling of points along the ray, which would significantly degrade performance. Instead, our focus is solely on determining which tangent-plane Gaussians the ray intersects on the unit sphere:
$\mathbf{x}_{\text{2D}}=\varphi_{\text{p}}\left(\left[\begin{matrix}\left(u-c_{x}\right)/f_{x}\\ \left(v-c_{y}\right)/f_{y}\\ 1\end{matrix}\right]\right)$ (16)
@@
where $c_{x}$, $c_{y}$, $f_{x}$, $f_{y}$ denote the intrinsic parameters of the pinhole camera. We then query the function values of these Gaussians for alpha blending to obtain the color
@@
$G_{\text{2D}}\left(\mathbf{x_{\text{2D}}}\right)=\exp\left\{-1/2\left(\mathbf{x_{\text{2D}}}-\varphi_{\text{p}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\right)^{\top}\left(\mathbf{J_{p}}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J_{p}}^{\top}\right)^{-1}\left(\mathbf{x_{\text{2D}}}-\varphi_{\text{p}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\right)\right\}\text{.}$ (17)
@@
Note that both $\mathbf{J_{p}}$ in Equation 14 and $\mathbf{J}$ in Equation 6 are matrices with a rank of 2. Consequently, the inverse of the covariance matrix of the 2D Gaussian is also obtained by skipping the third row and column. However, the third row and third column of $\mathbf{J_{p}}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J_{p}}^{\top}$ are not entirely zeros, as opposed to $\mathbf{J}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J}^{\top}$. To ensure the equation still holds, an invertible matrix $\mathbf{Q}$ needs to be used for left multiplication with $\mathbf{x_{\text{2D}}}$, $\varphi_{\text{p}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)$ and $\mathbf{J_{p}}$:
$\mathbf{Q}=\left[\begin{matrix}\frac{\mu_{z}}{\sqrt{\mu_{x}^{2}+\mu_{z}^{2}}}&0&-\frac{\mu_{x}}{\sqrt{\mu_{x}^{2}+\mu_{z}^{2}}}\\ -\frac{\mu_{x}\mu_{y}}{\sqrt{\mu_{x}^{2}+\mu_{z}^{2}}\sqrt{\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}}}&\frac{\sqrt{\mu_{x}^{2}+\mu_{z}^{2}}}{\sqrt{\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}}}&-\frac{\mu_{y}\mu_{z}}{\sqrt{\mu_{x}^{2}+\mu_{z}^{2}}\sqrt{\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}}}\\ \frac{\mu_{x}}{\sqrt{\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}}}&\frac{\mu_{y}}{\sqrt{\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}}}&\frac{\mu_{z}}{\sqrt{\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}}}\end{matrix}\right]\text{.}$ (18)
Since our projection does not rely on the perspective image plane $z=1$ as in the original 3D-GS [17], we can adapt to different camera models such as fisheye cameras and panoramic images by simply modifying the transformation from image space to camera space, as described in Equation 16.